We report on searches for B decays to charmless two-pseudoscalar and vector-pseudoscalar final states, using 9.7 × 10^{6}B\overline{B} pairs collected with the CLEO detector. Recent first observations include the decays B → π^{+}π^{−}, ρ^{0}π^{+}, and ρ^{+}π^{−}, where an average over charge conjugate modes is implied. Preliminary results for B → K^{∗+}π^{−} will also be presented. In addition, we discuss measurements of CP-violating asymmetries in the modes B → K^{±}π^{∓}, K^{±}π^{0}, K^{0}π^{±}, K^{±}η′, and ωπ^{±}.